OpenStudy (anonymous):

PLEASE HELP find the average rate of change of f(x) = 3x^2 + 1 (a) from 2 to 4 (b) from -2 to 0 and (c) from 2 to 5

OpenStudy (tkhunny):

Always, it is this: \(\dfrac{f(x_{2})-f(x_{1})}{x_{2}-x_{1}}\) You are given the x-values. Substitute and calculate!

OpenStudy (anonymous):

so 4-2 and divide? I am confused..

OpenStudy (tkhunny):

4-2 is the first denominator. You must also calculate the numerator, f(4) - f(2).

OpenStudy (anonymous):

so the denomiator is on the bottom?

OpenStudy (anonymous):

So how would I put it into an answer? 2,2?

OpenStudy (tkhunny):

You should know your numerator from your denominator. f(4) = 49 f(2) = 13 a) (49-13)/(4-2) -- Almost done.

OpenStudy (anonymous):

how did you get 49 and 13?

OpenStudy (anonymous):

I am sorry I am not good at this

OpenStudy (tkhunny):

You need to know how to use function notation. What does this mean? g(x) = 3x-4 If I ask for g(4), what does that mean? How does that notation work? I am quite worried about this. How dod you manage to get this close to calculus without encountering function notation? Very odd.

OpenStudy (anonymous):

plug 4 for x?

OpenStudy (tkhunny):

We do like to call it "substitute", but that's the idea. What do you get for g(4)?

OpenStudy (anonymous):

I don't know... I am sorry I am embarrassed

OpenStudy (tkhunny):

Yousaid yourself, sort of, substitute x = 4. If g(x) = 3x - 4, then g(4) = 3(4) - 4 = 12 - 4 = 8 g(9) = 3(9) - 4 = 27 - 4 = 23 Do you see how this works? You ARE going to have to review some algebra.